OpenStudy (anonymous):

What is the Derivative of f(x)=1/X? Would it be X^-1?

OpenStudy (anonymous):

-x^-2

OpenStudy (kinggeorge):

Write \[f(x)={1\over x} =x^{-1}\]Now just use the exponent rule. As Nameo0 just said, this is equal to \[-1 * x^{-2} = -{1\over x^2}\]

OpenStudy (anonymous):

Thank you ;-)

OpenStudy (anonymous):

Ok so then if I need to compelete a table I would use \[-x^{-2}\left( dx\right)\] to find dy? example if x=4 and dx = 1.000 then I have \[-4^{-2}\left( 1.000\right)\] =-.0625? and if my dx=delta(x) And use f(x+dx)-f(x) which would be \[4^{-1}+1.000-4^{-1}=1.000\] for delta y?